# The double gamma function and Vladimir Alekseevsky

### v2

@@ 3. Biography of Alekseevsky
31.10.1903. To theory of gammamorphic functions.
Clearly, the third talk corresponds to [Ale1904]. Sintsov mentions a disappeared article ‘On Riemann function $\zeta$’.
It seems clear that the topic of these works were the multiple gamma and related $\zeta$-functions of the type $\sum_{m,n\geqslant 0}(z+na+mb)^{-s}$ (something similar to Barnes [Bar1904]).
In any case, Alekseevsky has not published anything on this topic except claims in [Ale1889], [Ale1892], [Ale1894] quoted above [^18] Letter of Steklov to Lyapunov, 29.12.1902: Alekseevsky, as usual, drags on and does not submit his article [Yu. N.: the topic of the sentence is [Ale1904]], although he says that he will submit it ‘one of these days’, but this ‘one of these days’ has been going on since September..
3.5. The Tomsk Technological Institute, 1905–1916. The institute was founded in 1900 by chemist-technologist Efim Zubashev (about him, see [Gal2004]) and quickly became a center of education and science in Siberia. For instance, among professors of the beginning of the XX century there were Theodor Molien, one of the founders of representation theory of finite groups and theory of associative algebras [^19] See, e.g., Bourbaki [Bou1960], Section ‘Noncommutative algebra’, and selected works of Molien in [Mol1985]., and geologist Vladimir Obruchev.
According [Gran2003], Zubashev invited Alekseevsky to Tomsk in 1900 (the first year of work of the Institute). In April 1905, director Zubashev proposed to Alekseevsky the position of acting ordinary professor (full professor) of theoretical mechanics in Tomsk. Alekseevsky agreed and starting 28.11.1905 worked in Tomsk.
The paper [Gran2003] contains a lot of data on educational and administrative activity of Alekseevsky in Tomsk. We mention only a few formal details.
Students of the Tomsk Technological Institute were involved in the tumultuous events of the Russian Revolution 1905-07. Zubashev (who worked under pressure from both progressives and conservators) was accused of being loyal to the striking students and was exiled from Tomsk. Later, he returned to the director position but had to leave it ‘for health reasons’ (November 1907). The institute council elected Alekseevsky as the new director.
At the end of 1910 – 1911, in Russia there was a wave of student and professor protests against the reform of education initiated by the minister Lev Kasso. These protests involved the Institute. Several professors have been dismissed for oppositionism. Aleksevsky and the institute council tried to reinstate them. As a result, by the highest order, Alekseevsky left the director post [^20] The list of fired professors includes Molien and Obruchev. (8.06.1911). However, he continued his full professorship.
3.6. Publications of Alekseevsky. The list of scientific publications is contained in our bibliography, [Ale1888][Ale1904] (all of them belong to the Kharkov period of his work).
Russian library catalogues show several editions of his lecture courses and problem books (in the Kharkov Technological Institute and in the Tomsk Technological Institute):
@@
Also, [Sin1917] and [Gran2003] contains information about some additional lithographic editions of lecture courses in the Kharkov University and in the Tomsk Technological Institute.
@@
  • [Ale1884.1] Алексеевский В. П. Об интегрировании уравнения $\frac{d^{n}y}{dz^{n}}+\frac{\alpha}{z}\frac{d^{n-1}y}{dz^{n-1}}+\beta{y}=0,$ (Russian) Сообщ. и протоколы заседаний Матем. общ. при Император. Харьков. унив. 1884 года, 1884, 1, 41-64; Alexeiewsky V. P. On the integration of the equation $\frac{d^{n}y}{dz^{n}}+\frac{\alpha}{z}\frac{d^{n-1}y}{dz^{n-1}}+\beta{y}=0,$ Charkow Ges., 1884, 1, 41-64] JFM 16.0290.01212121Jahrbuch über die Fortschritte der Mathematik.
@@
  • [Ale1885] Алексеевский В. П., Об интегрировании одного линейного диффенциального уравнения $n$-го порядка (Russian) Сообщ. и протоколы заседаний Матем. общ. при Император. Харьков. унив. 1884 года, 1885, 3, 222-232.[ [Alexeiewsky V. P. On integration of one linear differential equation of $n$-th order, Charkow Ges., 1885, 3, 222-232] JFM 17.0328.01
  • [Ale1888] Алексеевский В. П. О функциях подобных функции гамма (Russian), Сообщ. Харьков. матем. общ. Вторая сер. [Charkow Ges.], 1:1 (1888), 169-192 [Alexeiewsky W. P. (Alexejewsky W.) On the functions that are similar to the gamma function.], Charkow Ges., 1:1 (1888), 169-192] JFM 22.0439.03.
  • [Ale1889] Алексеевский В. П. О функциях подобных функции гамма (Russian), Сообщ. Харьков. матем. общ. Вторая сер. [Charkow Ges.], 1:2 (1889), 193-238. [Alexeiewsky W. P. (Alexejewsky W.) On the functions that are similar to the gamma function., Charkow Ges., 1:2 (1889), 193-238] JFM 25.0751.02
@@
  • [Ale1895] Алексеевский В. П. Об автоморфной функции аналогичной экспонентной (Russian), Сообщ. Харьков. матем. общ. Вторая сер. [Charkow Ges.], 4 (1895), 253-262 [Alexeiewsky V. P. On an automorphic function that is analogous to the exponential, Charkow Ges., 4 (1895), 253-262] JFM 26.0466.02
@@
  • [Fock1997] Fock V. V. Dual Teichmüller spaces. Preprint https://arxiv.org/pdf/dg-ga/9702018v1.pdf
@@
  • [Gal2004] Galanova R. A. His thoughts and deeds were devoted to Siberia: about the first rector of TTI (TPU) [^22] The Tomsk Techological Institute (TTI) changed couple of names, since 1991 it is The Tomsk Technological University, (TTU) E. L. Zubashev. (Russian). Tomsky polytehnik, 2004, 10, 160-163.
@@
  • [Gran2003] Granina Yu. G. The professor-mathematician, the second director of TTI V. P. Alekseevsky (Russian), Izvestiya Tomskogo polytechnicheskogo universiteta, [Bulletin of the Tomsk polytechnical university], 306, no. 3., (2003) 162-167.
@@
  • [KarKu2024] Karp D., Kuznetsov A. Extending the Meijer G-function. Preprint https://arxiv.org/abs/2403.05708.
@@
  • [Kru1967.1] Krulikovskiĭ. N. N. Mathematics in the Tomsk Technological Institute. (Russian) In History of national mathematics in four volumes. Vol. 2, 1801-1917. (eds. Z. Shtokalo, A. N. Bogoljubov, Ju. A. Mitropol’skiĭ, I. B. Pogrebysskiĭ, E. Ja. Remez, K. A. Rybnikov, Ju. D. Sokolov and V. S. Sologub), 524-527. Naukova Dumka, Kiev, 1967.
  • [Mari1978] Маричев О.И. Метод вычисления интегралов от специальных функций (теория и таблицы формул) (Russian), Наука и техника, Минск, 1978 г. [Marichev O. I The method of evaluation of integrals of special functions (theory and tables of formulas), Nauka i Tehnika, Minsk, 1978]; English version Marichev O. I. Handbook of integral transforms of higher transcendental functions: theory and algorithmic tables. John Wiley, New York, 1983.
@@
  • [Mol1985] Molin F. È. [Molien Th.] Number systems. (Russian) Translated from the German by L. A. Bokut’, N. N. Krulikovskiĭ. and I. V. L’vov. Edited by A. I. Kostrikin. With appendices by Krulikovskiĭ, Bokut’. and L’vov. Nauka, Sibirsk. Otdel., Novosibirsk, 1985. 126 pp. Zbl 0607.01025
@@
  • [Ste1904] Stekloff W. A. Remarques relatives aux formules sommatoires d’Euler et de Bool (French), Сообщ. Харьков. матем. общ. Вторая сер. [Charkow Ges.], 8 (1904), 136-195. JFM 35.0263.01
@@
Institute for Information Transmission Problems;
Moscow State University, MechMath Faculty

### v3

@@ 5. Alekseevsky: his historiography, lost work, and publications
31.10.1903. Towards the theory of gammamorphic functions.
Clearly, the third talk corresponds to [Ale1904].
Sintsov [Sin1917] mentions that an article ‘On Riemann function $\zeta$’ disappeared and writes that there is hope to find it in Alekseevsky’s posthumous papers.
It seems clear that the topic of these works were the multiple gamma functions and related $\zeta$-functions of the type
$\sum_{m,n\geqslant 0}(z+n\omega_{1}+m\omega_{2})^{-s}$
(something similar to Barnes [Bar1904]).
In any case, Alekseevsky has not published anything on this topic except the claims in [Ale1889], [Ale1892], [Ale1894] quoted above [^23] Letter of Steklov to Lyapunov, 29.12.1902: Alekseevsky, as usual, drags on and does not submit his article [Yu. N.: clearly, this sentence refers to [Ale1904]], although he says that he will submit it ‘one of these days’, but this ‘one of these days’ has been going on since September..
5.3. Alekseevsky’s publications. The list of his research publications is contained in our bibliography, [Ale1884.1][Ale1904] (all of them belong to the Kharkov period of his work).
Russian library catalogues show several editions of his lecture courses and problem books (in Kharkov Technological Institute and in Tomsk Technological Institute):
@@
Also, [Sin1917] and [Gran2003] contains information about some additional lithographic editions of his lecture courses in Kharkov University and in Tomsk Technological Institute.
6. Conclusion
The double gamma was introduced in the work of Vladimir Alekseevsky in 1888-1889, he also derived main identities involving this function (cf. the list of identities [AlKu2023]). His publication [Ale1894] had a big influence on the early works by Ernest Barnes, who also introduced and investigated the multiple gamma function [Bar1904]. The double gamma function was Alekseevsky’s main work and he was forgotten for the same period that the double gamma itself was forgotten.
In this paper, we observed an interesting case: that the results of a work, which were very difficult to replicate, were forgotten for 70 years. This happened despite the fact that some of the authors writing about the double gamma, were well-known and that the results were vivid. The cause was probably an unfortunate confluence of several circumstances.
Modern mathematicians use complicated languages, having trouble understanding each other. Large losses of nontrivial information over the intervals of several decades are unavoidable (and is, in fact, already occuring). In our case, however, the papers on special functions and complex analysis of 1880-1910s were written in a simple language one accessible for mathematicians of 1970-s and since. They could have been easily read and used.
@@ 6. Conclusion
  • [Ale1884.1] Алексеевский В. П. Об интегрировании уравнения $\frac{d^{n}y}{dz^{n}}+\frac{\alpha}{z}\frac{d^{n-1}y}{dz^{n-1}}+\beta{y}=0,$ (Russian) Сообщ. и протоколы заседаний Матем. общ. при Император. Харьков. унив. 1884 года, 1884, 1, 41-64; [Alexeiewsky V. P. On the integration of the equation $\frac{d^{n}y}{dz^{n}}+\frac{\alpha}{z}\frac{d^{n-1}y}{dz^{n-1}}+\beta{y}=0,$ Charkow Ges., 1884, 1, 41-64] JFM [^24] JFM database; the electronic version of Jahrbuch über die Fortschritte der Mathematik, 1868-1942. It is contained in the Zentralblatt für Mathematik database. 16.0290.01.
@@
  • [Ale1885] Алексеевский В. П., Об интегрировании одного линейного диффенциального уравнения nn-го порядка (Russian) Сообщ. и протоколы заседаний Матем. общ. при Император. Харьков. унив. 1884 года, 1885, 3, 222-232. [Alexeiewsky V. P. On integration of one linear differential equation of $n$-th order, Charkow Ges., 1885, 3, 222-232] JFM 17.0328.01
  • [Ale1888] Алексеевский В. П. О функциях подобных функции гамма (Russian), Сообщ. Харьков. матем. общ. Вторая сер., 1:1 (1888), 169-192 [Alexeiewsky W. P. (Alexejewsky W.) On the functions that are similar to the gamma function.], Charkow Ges., 1:1 (1888), 169-192] JFM 22.0439.03.
  • [Ale1889] Алексеевский В. П. О функциях подобных функции гамма (Russian), Сообщ. Харьков. матем. общ. Вторая сер., 1:2 (1889), 193-238. [Alexeiewsky W. P. (Alexejewsky W.) On the functions that are similar to the gamma function., Charkow Ges., 1:2 (1889), 193-238] JFM 25.0751.02
@@
  • [Ale1895] Алексеевский В. П. Об автоморфной функции аналогичной экспонентной (Russian), Сообщ. Харьков. матем. общ. Вторая сер., 4 (1895), 253-262 [Alexeiewsky V. P. On an automorphic function that is analogous to the exponential, Charkow Ges., 4 (1895), 253-262] JFM 26.0466.02
@@
  • [Fock1997] Fock V. V. Dual Teichmüller spaces. Preprint
    . https://arxiv.org/pdf/dg-ga/9702018v1.pdf
@@
  • [Gal2004] Galanova R. A. His thoughts and deeds were devoted to Siberia: about the first rector of TTI (TPU) [^25] Tomsk Techological Institute (TTI) changed couple of names, since 1991 it is Tomsk Technological University, (TTU). E. L. Zubashev. (Russian). Tomsky polytehnik, 2004, 10, 160-163.
  • [Gal2009] Galanova R. A. Efim Luk’yanovich Zubashev: a biographical essay. (Russian) Tomsk Polytechnical University, Tomsk, 2009.
@@
  • [Gran2003] Granina Yu. G. The professor-mathematician, the second director of TTI V. P. Alekseevsky (Russian), Izvestiya Tomskogo polytechnicheskogo universiteta, [Bulletin of Tomsk polytechnical university], 306, no. 3., (2003) 162-167.
@@
  • [Gus1953] Gussov, V. V. Development of the theory of cylinder functions in Russia and the USSR. (Russian) Istor.-Mat. Issled. (1953), 355-475.
@@
  • [KarKu2024] Karp D., Kuznetsov A. Extending the Meijer G-function. Preprint
    . https://arxiv.org/abs/2403.05708.
@@
  • [KashN2011] Kashaev R. M., Nakanishi T. Classical and quantum dilogarithm identities. SIGMA Symmetry Integrability Geom. Methods Appl. 7 (2011), Paper 102, 29 pp.
@@
  • [Kru1967.1] Krulikovskiĭ. N. N. Mathematics in Tomsk Technological Institute. (Russian) In History of national mathematics in four volumes. Vol. 2, 1801-1917. (eds. Z. Shtokalo, A. N. Bogoljubov, Ju. A. Mitropol’skiĭ, I. B. Pogrebysskiĭ, E. Ja. Remez, K. A. Rybnikov, Ju. D. Sokolov and V. S. Sologub), 524-527. Naukova Dumka, Kiev, 1967.
  • [Mari1978] Marichev O. I The method of evaluation of integrals of special functions (theory and tables of formulas) (Russian), Nauka i Tehnika, Minsk, 1978; English version: Marichev O. I. Handbook of integral transforms of higher transcendental functions: theory and algorithmic tables. John Wiley, New York, 1983.
@@
  • [Mark1981] Markushevich A. I. Theory of analytic functions. (Russian). In Mathematics of the XIXth century, Geometry, Theory of analytic functions (eds. A. N. Kolmogorov, A. P. Yushkevich), Moscow 1981, 115-269 (1981); English transl.: Laptev B. L., Rozenfeld B. A., Markushevich A. I. Mathematics of the 19th century. Geometry, analytic function theory. Birkhäuser, Basel, 1996.
  • [Mol1985] Molin F. È. [Molien Th.] Number systems. (Russian) Translated from the German by L. A. Bokut’, N. N. Krulikovskiĭ. and I. V. L’vov. Edited by A. I. Kostrikin. With appendices by Krulikovskiĭ, Bokut’. and L’vov. Nauka, Sibirsk. Otdel., Novosibirsk, 1985. 126 pp. Zbl [^27] The database Zentralblatt für Mathematik. 0607.01025
@@
  • [PBM1989] Prudnikov A. P., Brychkov Yu. A., Marichev O. I. Calculation of integrals and the Mellin transform. (Russian) Mathematical analysis, Vol. 27, 3-146, Progress in Science and Technology, VINITI, 1989; Translated in J. Soviet Math. 54 (1991), no. 6, 1239-1341.
@@
  • [Rim1957.1] Rimskiĭ-Korsakov B. S. An investigation on theory of generalizations of the Gamma-function (Russian). Dissertation for the degree of candidate of physical and mathematical sciences, Moscow, 1957.
@@
  • [Scho1884] Schobloch, J. A. Über Beta- und Gammafunctionen. (German) Halle. Nebert (1884). JFM 16.0395.01.
@@
  • [Srin2007] Srinivasan G. K. The gamma function: an eclectic tour. Amer. Math. Monthly 114 (2007), no. 4, 297-315.
  • [Ste1904] Stekloff W. A. [Steklov V. A.] Remarques relatives aux formules sommatoires d’Euler et de Bool (French), Сообщ. Харьков. матем. общ. Вторая сер. [Charkow Ges.], 8 (1904), 136-195. JFM 35.0263.01
@@
  • [Wae1985] van der Waerden B. L. A history of algebra. From al-Khwārizmī. to Emmy Noether. Springer-Verlag, Berlin, 1985.
@@
  • [Wor2000] Woronowicz S.L. Quantum exponential function, Rev. Math. Phys. 12 (2000), 873-920.
  • [Yush1968] Yushkevich A. P. History of mathematics in Russia up to 1917. (Russian) Nauka, Moscow, 1968, 591 pp.
University of Graz;
High School of Modern Mathematics MIPT;
Moscow State University, MechMath Faculty;
Univesity of Vienna, Faculty of Mathematics.